OpenStudy (anonymous):

Hey! So I really need help with my trig homework, and yes I have tried for the life of me to understand it but I just don't!!! I need to determine if the sides of these triangles are acute, obtuse, or a right triangle and I don't get it!!! Help!! 8. Sides have lengths 6 in., 7 in., and 8 in. 9. Sides have lengths 36 ft., 39 ft., and 15 ft. 10. Sides have lengths 9 m, 33 m and 41 m Please help!! And I'm not trying to cheat at all.

OpenStudy (asad786):

ok whats wrong

OpenStudy (rational):

`Sides have lengths 6 in., 7 in., and 8 in.1` take the longest side and square it : 8^2 = 64 take the sum of squares of smaller sides : 6^2 + 7^2 = 85 since 64 is less than 85, the triangle is ACUTE

OpenStudy (anonymous):

Thank you!!

OpenStudy (rational):

In general, a triangle with sides : "a, b, c" is acute if `c^2 < a^2 + b^2` where "c" is the longest side

OpenStudy (rational):

it is obtuse if `c^2 > a^2 + b^2` it is right if `c^2 = a^2 + b^2`

OpenStudy (rational):

see if u can work next problem on ur own

OpenStudy (anonymous):

Okay so it would be: 39^2= 1521 36^2+15^2=1521 so it would be a right triangle. Right?

OpenStudy (anonymous):

I think I've got it!!! Thank you!!!!!!!

OpenStudy (rational):

Yes! good job!